OpenStudy (anonymous):

USE IMPLICIT X^3-2XY+Y^2=4 (1,-1)

OpenStudy (anonymous):

3x^2 - 2x(dy/dx) - 2y + 2y(dy/dx) = 0 (dy/dx)(-2x + 2y) = 2y - 3x^2 dy/dx = (2y - 3x^2)/(2y - 2x), then plug in dy/dx = -5/-4 = 5/4 You can plug in the values before isolating the derivative on one side and it would be quicker, but oh well.

OpenStudy (anonymous):

SO PLUG IN (1,-1) in (2y-3x^2)/(2y-2x)

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

thank you so much